OpenStudy (anonymous):

If vector A⃗ has components Ax and Ay and makes an angle θ with the +x axis, then...

OpenStudy (anonymous):

tanθ =Ax/Ay ?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

isnt it supposed to be y/x instead?

OpenStudy (anonymous):

wait crap. Yea. It's y/x.

OpenStudy (anonymous):

A=Ax+Ay (where Ais the magnitude of A⃗ ) θ =AyAx cosθ = AxA2x+A2y√ tanθ =AxAy which of these is true? the first one?

OpenStudy (anonymous):

Hmmm not quite sure...might want to look that up or ask someone else. I don't want to give any false info.

OpenStudy (anonymous):

k cool thanks

OpenStudy (isaiah.feynman):

yes the tangent of the angle is ay/ax

OpenStudy (anonymous):

The vector is F = (Ax,Ay) and it has direction @ such that tan@ = Ay/Ax